OpenStudy (anonymous):

I really need help!!!!!!!!!!!:( Solve for x: 8x2 + 64x = 0

OpenStudy (anonymous):

i will give medals

OpenStudy (anonymous):

Calculate it! 8x(2)+64x=0 I will try to solve your equation. 80x=0(Simplify both sides of the equation) (80x) over 80 = 0 over 80 (Divide both sides by 80) x=0

OpenStudy (anonymous):

Please give medals and fan me if this helped! :D

OpenStudy (anonymous):

I got x=0 ad x=-8 is that right? because you just gave me the part of the answer ??

OpenStudy (anonymous):

Yes, x does = 0

OpenStudy (anonymous):

what about x=-8 did I get that part right?

OpenStudy (anonymous):

But not -8.

OpenStudy (anonymous):

You can try it

OpenStudy (anonymous):

why not -8

OpenStudy (anonymous):

Is the 8x2 supposed to be an 8x^2?

OpenStudy (anonymous):

Now, if the equation was: 8x2 + 64x = then it would be =80x

OpenStudy (anonymous):

i took a test and it said that the answer was x=0 and x=-8 i am just trying to figure out why

OpenStudy (anonymous):

Hmm.....I don't understand why it says that. Do you have a copy by any chance?

OpenStudy (anonymous):

yeah here was the question and options Solve for x: 8x2 + 64x = 0 (2 points) x = 0, x = -8 x = 0, x = 8 x = -8, x = 0, x = 8 x = 8, x = 64

OpenStudy (anonymous):

Like a screen capture of link

OpenStudy (anonymous):

Try -8.why did you make it positive?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Do you do Laurel Springs?

OpenStudy (anonymous):

That is weird.......maybe the class teaches it differently

OpenStudy (anonymous):

I do Florida virtual

OpenStudy (anonymous):

Oh. I am in Florida and do online school! :D It seemed familiar but I do Keystone National Online School

OpenStudy (anonymous):

so what do you think?

OpenStudy (anonymous):

acacia, how many ways of solving a quadratic do you know?

OpenStudy (anonymous):

Sorry I can only get x=0 :(

OpenStudy (anonymous):

well i do this program but i live in Oregon

OpenStudy (anonymous):

umm 4 ways i think

OpenStudy (anonymous):

Which one do you think would be the best for this equation?

OpenStudy (anonymous):

the one where you use this formula x=−b +/− √b2−4ac /2a

OpenStudy (anonymous):

Okay, so in this case a=8,b=64,and c=0 you can substitute these into the equation and solve for x

OpenStudy (anonymous):

ok I got it thanks so much i will give you a medal and a fan!

OpenStudy (anonymous):

You're welcome. Would you like to see the other ways that might have worked?

OpenStudy (anonymous):

no I think I got it thanks thou by!!